### v1

# Asymptotically normal estimators in high-dimensional linear regression

Abstract
We establish asymptotic normality for estimators in high‑dimensional linear regression by proving weak convergence in a separable Hilbert space, thereby enabling direct use of standard asymptotic tools, for example, the continuous mapping theorem. The approach allows the number of non‑zero coefficients to grow, provided only a fixed number have moderate magnitude. As an application, we test linear hypotheses with a statistic whose null limit is a finite weighted sum of independent chi‑squared variables, yielding plug‑in critical values with asymptotically correct size.
keywords: high-dimensional inference; linear regression; sparsity; weak convergence
1 Introduction
We develop a new framework for establishing the asymptotic normality of estimators of high-dimensional parameters via weak convergence in the Hilbert space $\ell^{2}$, thereby enabling the use of standard asymptotic tools for broad classes of continuous functionals. Existing results typically assume fixed sparsity, namely that the number of non-zero coefficients does not grow, or restrict attention to low-dimensional linear functionals. We relax these restrictions and allow the sparsity level to diverge; in particular, the number of non-zero coefficients of small magnitude may increase with the sample size. For clarity of exposition, we focus on linear regression with independent and identically distributed (i.i.d.) sub-Gaussian covariates and errors.
Numerous methods for high-dimensional sparse estimation have been proposed, including the least absolute shrinkage and selection operator (lasso) (Tibshirani, 1996), the adaptive lasso (Zou, 2006), the Dantzig selector (Candes and Tao, 2007), and losses with nonconvex penalties such as the smoothly clipped absolute deviation and the minimax concave penalty (Fan and Li, 2001; Zhang, 2010). Foundational results include oracle inequalities and support-recovery guarantees (Bickel et al., 2009; Wainwright, 2009). On the inferential side, post-selection procedures have been analyzed (Belloni and Chernozhukov, 2013); separately, two-step estimators for stochastic process models have been developed (Fujimori and Tsukuda, 2026); and de-biasing methods yield asymptotically normal estimators of low-dimensional functionals (Zhang and Zhang, 2014; van de Geer et al., 2014; Javanmard and Montanari, 2014).
Accordingly, an extensive literature has developed non-asymptotic, high-dimensional central limit theorems and Gaussian approximations under metrics such as the Kolmogorov and Wasserstein distances. Representative results include Gaussian approximations for maxima of sums of centred independent random vectors (Chernozhukov et al., 2013, 2017), Stein kernel–based approximations accommodating exchangeable pairs and certain nonlinear statistics (Fang and Koike, 2021), rate refinements (Koike, 2023), and extensions to convex polytopes and degenerate cases (Fang et al., 2023).
Rather than deriving metric‑specific non-asymptotic bounds, we establish weak convergence of a high-dimensional estimator in $\ell^{2}$, bringing standard asymptotic tools such as the continuous mapping theorem to bear on a broad class of continuous functionals. Coordinatewise normality and normality of fixed‑dimensional linear functionals follow immediately. We illustrate the framework by constructing a test of linear hypotheses, contrasting it with projection‑based procedures that accommodate dense nuisance parameters (Zhu and Bradic, 2018) and hypothesis‑adaptive tests designed for high power (Zhang et al., 2025); our approach handles non-sparsity via a different route.
Throughout the paper, we use the following notation. We write $k]=\{1,\ldots,k\}$ for $k\in\mathbb{N}$. For $q\in\mathbb{N}\cup\{\infty\}$, write $\|\bm{v}\|_{q}$ for the $\ell^{q}$ norm: for $\bm{v}=(v_{1},\ldots,v_{p})^{\top}\in\mathbb{R}^{p}$ with $p\in\mathbb{N}$, $\|\bm{v}\|_{q}=(\sum_{j=1}^{p}|v_{j}|^{q})^{1/q}$ $(q<\infty)$, $\|\bm{v}\|_{\infty}=\max_{1\leq j\leq p}|v_{j}|$. For an infinite sequence $\bm{v}=(v_{j})_{j\geq 1}\in\ell^{q}$, the same formulae apply. For a bounded linear operator $\bm{A}:\ell^{2}\to\ell^{2}$ or a matrix $\bm{A}$ in finite dimensions, let $\|\bm{A}\|_{\mathrm{op}}=\sup_{\|\bm{x}\|_{2}=1}\|\bm{A}\bm{x}\|_{2}$ denote the $\ell^{2}$‑induced operator norm. For a vector $\bm{v}=(v_{j})_{j\geq 1}$ and an index set $T\subset\mathbb{N}$ with $|T|<\infty$, write $\bm{v}_{T}=(v_{j})_{j\in T}$, which is viewed as an element of $\mathbb{R}^{|T|}$. In the finite‑dimensional case $\bm{v}\in\mathbb{R}^{p}$, we restrict to $T\subset[p]$. Similarly, for a $p\times p$ matrix $\bm{A}=(A_{ij})_{i,j\in[p]}$ and index sets $T,T^{\prime}\subset[p]$, write $\bm{A}_{T,T^{\prime}}=(A_{ij})_{i\in T,j\in T^{\prime}}$. For a random variable $X$, we write the sub-Gaussian norm of $X$, if it exists, as $\|X\|_{\psi_{2}}$, that is, $\|X\|_{\psi_{2}}=\inf\{C>0:\,\mathsf{E}[\exp({X^{2}}/{C^{2}})]\leq 2\}$. For a $d$-dimensional random vector $\bm{X}$, the sub-Gaussian norm of $\bm{X}$ is $\|\bm{X}\|_{\psi_{2}}=\sup_{\|\bm{u}\|_{2}=1}\|\langle\bm{u},\bm{X}\rangle\|_{\psi_{2}}$. For $d\in\mathbb{N}$, $\bm{\mu}\in\mathbb{R}^{d}$, and $\bm{V}\in\mathbb{R}^{d\times d}$ positive semi‑definite, we write $\mathcal{N}_{d}(\bm{\mu},\bm{V})$ for the $d$-dimensional normal distribution with mean $\bm{\mu}$ and covariance $\bm{V}$, with the subscript omitted when $d=1$. For random elements $X$ and $Y$ and a sequence $\{X_{n}\}_{n\in\mathbb{N}}$ of random elements, we write $X\stackrel{{\scriptstyle d}}{{=}}Y$ if $X$ and $Y$ have identical distributions, and we write $X_{n}\to^{d}X$ and $X_{n}\to^{p}X$ for convergence in distribution and in probability as $n\to\infty$, respectively.
2 Model set‑up and main result
For each sample size $n\in\mathbb{N}$, consider $(p_{n}+1)$-dimensional observations $\{(Y_{i,n},\bm{Z}_{i,n})\}_{i\in[n]}$ satisfying
$Y_{i,n}=\bm{\theta}_{n}^{\top}\bm{Z}_{i,n}+\epsilon_{i,n},\quad i\in[n],$
where $\bm{\theta}_{n}\in\mathbb{R}^{p_{n}}$ and $p_{n}$ may diverge with $n$. The sequence $\{(\bm{Z}_{i,n},\epsilon_{i,n})\}_{i\in[n]}$ are i.i.d. with $\mathsf{E}(\bm{Z}_{1,n})=\bm{0}$, $\mathsf{E}(\epsilon_{1,n})=0$, $\mathsf{V}(\epsilon_{1,n})=\sigma_{n}^{2}>0$, and $\bm{Z}_{1,n}$ independent of $\epsilon_{1,n}$. Let $\bm{\theta}_{0,n}$ denote the true parameter vector. Hereafter, we suppress the subscript $n$ when no confusion can arise. Write $\bm{\theta}_{0}=(\theta_{01},\ldots,\theta_{0p})^{\top}$. Define the index sets
$T_{0}:=\{j\in[p]:\theta_{0j}\neq 0\},\quad T_{\gamma}:=\{j\in[p]:|\theta_{0j}|>\gamma\},\quad T_{W}:=\{j\in[p]:\theta_{\min}\leq|\theta_{0j}|\leq\eta_{p}\},$
where $\theta_{\min}:=\min_{j\in T_{0}}|\theta_{0j}|$, $\gamma>0$ is a fixed constant independent of $n$ and $p$, and $\eta_{p}$ is a positive sequence depending on $p$. Let $s_{0}:=|T_{0}|$ and $s_{\gamma}:=|T_{\gamma}|$ denote the corresponding support sizes, where $s_{0}$ may diverge as $n\to\infty$. The next assumption specifies the conditions on $\gamma$, $s_{0}$, $s_{\gamma}$, and $\eta_{p}$.
Assumption 1.
The following conditions hold:
  1. (i)
    $s_{\gamma}$ is fixed and independent of $n$ and $p$.
  2. (ii)
    $T_{0}\setminus T_{\gamma}=T_{W}$.
  3. (iii)
    $ns_{0}\eta_{p}^{2}\to 0$ as $n\to\infty$.
  4. (iv)
    there exists $\theta_{\max}>0$, independent of $n$ and $p$, such that $\|\bm{\theta}_{0}\|_{\infty}\leq\theta_{\max}$.
Remark 1.
Assumption 1-(ii) states that the non-zero coefficients outside $T_{\gamma}$ are precisely the weak signals, and Assumption 1-(iii) requires that their aggregate magnitude is asymptotically negligible. Under Assumption 1, $(\bm{\theta}_{0,n},0,0,\ldots)\in\mathbb{R}^{\infty}$ lies in $\ell^{2}$ uniformly in $n$, because
$\|\bm{\theta}_{0}\|_{2}^{2}=\sum_{j\in T_{\gamma}}|\theta_{0j}|^{2}+\sum_{j\in T_{0}\setminus T_{\gamma}}|\theta_{0j}|^{2}\leq\sum_{j\in T_{\gamma}}|\theta_{0j}|^{2}+\eta_{p}^{2}s_{0}\leq s_{\gamma}\theta_{\max}^{2}+o(1)=O(1)$
as $n\to\infty$.
Let $\hat{\bm{\theta}}_{n}$ be a sparse estimator for $\bm{\theta}$ such that $\mathsf{P}(\|\hat{\bm{\theta}}_{n}-\bm{\theta}_{0}\|_{\infty}>c_{n})\to 0$ as $n\to\infty$, where the rate $c_{n}$ satisfies that $c_{n}=o(1)$ as $n\to\infty$. Given a threshold sequence $\{\tau_{n}\}_{n\in\mathbb{N}}$, define the sequence $\{\hat{T}_{n}\}_{n\in\mathbb{N}}$ of an estimator of $T_{0}$ by
$\hat{T}_{n}=\{j\in[p]:|\hat{\theta}_{nj}|>\tau_{n}\}.$
Remark 2.
For example, the lasso estimator and the Dantzig selector satisfy
$\|\hat{\bm{\theta}}_{n}-\bm{\theta}_{0}\|_{\infty}=O_{p}(n^{-1/2}(\log p)^{1/2})$
as $n\to\infty$. Accordingly, in such cases we may take $c_{n}\asymp n^{-1/2}(\log p)^{1/2}$.
The following condition suffices to ensure selection consistency.
Assumption 2.
The following condition holds: $c_{n}+\eta_{p}<\tau_{n}<\gamma-c_{n}$.
We have the following selection consistency.
Lemma 2.1.
Let Assumptions 12 hold. Then, $\mathsf{P}(\hat{T}_{n}=T_{\gamma})\to 1$ as $n\to\infty$.
We define the ordinary least squares estimator $\tilde{\bm{\theta}}_{n}=(\tilde{\bm{\theta}}_{n1},\ldots,\tilde{\bm{\theta}}_{np})$ after variable selection as the solution to the estimating equation
$n^{-1}\sum_{i=1}^{n}\bm{Z}_{i\hat{T}_{n}}(Y_{i}-\bm{\theta}_{\hat{T}_{n}}^{\top}\bm{Z}_{i\hat{T}_{n}})=\bm{0},\quad\bm{\theta}_{\hat{T}_{n}^{c}}=\bm{0}.$
[…]
where $\{\bm{e}_{j}\}_{j\in\mathbb{N}}$ is the canonical basis of $\mathbb{R}^{\infty}$. Define
$\bm{J}_{n}:=n^{-1}\sum_{i=1}^{n}\bm{Z}_{i}\bm{Z}_{i}^{\top},\quad\bm{J}:=\,\mathsf{E}[\bm{Z}_{1}\bm{Z}_{1}^{\top}].$ (2.1)
In addition, we impose the following conditions.
Assumption 3.
The following conditions hold:
  1. (i)
    There exists $\lambda_{*}>0$ and $\Lambda_{*}>0$, independent of $n$ and $p$, such that $\lambda_{*}\leq\Lambda_{\min}(\bm{J}_{T_{\gamma},T_{\gamma}})\leq\Lambda_{\max}(\bm{J}_{T_{\gamma},T_{\gamma}})\leq\Lambda_{*}$.
  2. (ii)
    $\bm{Z}_{1},\ldots,\bm{Z}_{n}$ are i.i.d. copies of a sub-Gaussian random vector $\bm{Z}$.
  3. (iii)
    $\epsilon_{1},\ldots,\epsilon_{n}$ are i.i.d. copies of a sub-Gaussian random variable $\epsilon$.
Under Assumption 3-(ii), there exist positive constants $K_{0}$ and $K_{p}$ such that $\|Z_{j}\|_{\psi_{2}}\leq K_{0}$ for $j\in[p]$ and $\|\bm{Z}\|_{\psi_{2}}\leq K_{p}$, where $K_{0}$ is independent of $n$ and $p$, while $K_{p}$ may depend on $p$. Under Assumption 3-(iii), there exists a positive constant $K_{\epsilon}$ such that $\|\epsilon\|_{\psi_{2}}\leq K_{\epsilon}$, where $K_{\epsilon}>0$ is independent of $n$, $p$, and $s_{0}$. Hereafter, we use these constants $K_{0}$, $K_{p}$, and $K_{\epsilon}$.
Remark 3.
Under Assumption 3-(ii), there exists a universal constant $C>0$ such that $\|\bm{Z}_{S}\|_{\psi_{2}}\leq CK_{0}|S|^{1/2}$ for any $S\subset[p]$ whose cardinality is fixed and independent of $n$ and $p$.
Under the assumption that $s_{\gamma}$ is independent of $p$, the following lemma holds.
Lemma 2.2.
Let Assumptions 1 and 3 hold. Then, $\|\bm{J}_{nT_{\gamma},T_{\gamma}}-\bm{J}_{T_{\gamma},T_{\gamma}}\|_{\mathrm{op}}=o_{p}(1)$ as $n\to\infty$.
We now state the main limit theorem.
Theorem 2.3.
[…]
Remark 4.
Fujimori and Tsukuda (2026) analyze asymptotic normality for high‑dimensional estimators via weak convergence in $\ell^{2}$, including settings with a possibly infinite‑dimensional nuisance parameter and dependent observations. Their analysis is confined to the fixed-sparsity regime, which materially streamlines the analysis. In contrast, we allow the support size $s_{0}\to\infty$ while keeping the number $s_{\gamma}$ of coefficients of moderate magnitude fixed.
To prove Theorem 2.3, it suffices to establish Lemma 2.4; see, for example, Chapter 1.8 of van der Vaart and Wellner (2023).
Lemma 2.4.
Let Assumptions 13 hold.
  1. (i)
    For every $\eta>0$, $\mathsf{P}(\|n^{1/2}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}^{c}}\|_{2}^{2}>\eta)\to 0$ as $n\to\infty$.
  2. (ii)
    For every $\bm{u}\in\ell^{2}$, $𝓡\langle\bm{u},\bm{\mathcal{R}}_{n}\rangle\to^{d}\langle\bm{u},\bm{\mathcal{R}}\rangle$ as $n\to\infty$.
3 Application: testing linear hypotheses
As an application of Theorem 2.3, we consider testing a linear hypothesis for the high-dimensional parameter $\bm{\theta}$. Let $q(\leq p)$ be an integer, $\bm{A}\in\mathbb{R}^{q\times p}$ a given matrix, $\bm{b}\in\mathbb{R}^{q}$ a given vector. Let $\bm{\Sigma}_{A}=\bm{A}_{[q],T_{\gamma}}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{A}_{[q],T_{\gamma}}^{\top}$, where $\bm{J}$ is defined in (2.1). Moreover, let $r$ be the rank of $\bm{\Sigma}_{A}$, and let $\Lambda_{1}\leq\ldots\leq\Lambda_{r}$ be the non-zero eigenvalues of $\bm{\Sigma}_{A}$. Since $s_{\gamma}$ is independent of $n$ and $p$, we have $r\leq\min(s_{\gamma},q)$, so that $r$ is fixed. To test $\mathcal{H}_{0}:\bm{A}\bm{\theta}=\bm{b}$ against $\mathcal{H}_{1}:\bm{A}\bm{\theta}\neq\bm{b}$, consider
$W_{n}:=n\|\bm{A}\tilde{\bm{\theta}}_{n}-\bm{b}\|_{2}^{2}.$
We now state the condition on $q$.
Assumption 4.
The following condition holds: $q$ is fixed and independent of $n$ and $p$.
Remark 5.
An extension to the case $q=q_{n}\to\infty$ is also possible under an additional assumption that $\bm{A}$ defines a bounded linear operator on $\ell^{2}$.
The asymptotic null distribution of $W_{n}$ is given in the following corollary.
Corollary 3.1.
Let Assumptions 14 hold. Under $\mathcal{H}_{0}$,
$W_{n}\to^{d}\sigma^{2}\sum_{j=1}^{r}\Lambda_{j}\chi_{1,j}^{2}$
as $n\to\infty$, where $\chi_{1,1}^{2},\ldots,\chi_{1,r}^{2}$ are independent and $\chi^{2}(1)$-distributed random variables.
We consider the following simple estimators of $\sigma^{2}$ and $\Lambda_{1},\ldots,\Lambda_{r}$. For $\sigma^{2}$, define
$\hat{\sigma}_{n}^{2}=n^{-1}\sum_{i=1}^{n}\left(Y_{i}-\tilde{\bm{\theta}}_{n}^{\top}\bm{Z}_{i}\right)^{2}.$
For $\Lambda_{j}$ ($j\in[r]$), define $\hat{\Lambda}_{n,j}$ as the $j$th largest eigenvalue of
$\hat{\bm{\Sigma}}_{nA}=\bm{A}_{[q],T_{\gamma}}\bm{J}_{nT_{\gamma},T_{\gamma}}^{-1}\bm{A}_{[q],T_{\gamma}}^{\top}.$
The following lemma establishes consistency of these estimators.
Lemma 3.2.
Let Assumptions 14 hold. Then, $\hat{\sigma}_{n}^{2}\to^{p}\sigma_{0}^{2}$ and $\max_{j\in[r]}|\hat{\Lambda}_{n,j}-\Lambda_{j}|\to^{p}0$ as $n\to\infty$.
Let $c_{\alpha}$ and $\hat{c}_{\alpha}$ denote the upper $(1-\alpha)$-quantiles of the distributions of
$\sigma^{2}\sum_{j=1}^{r}\Lambda_{j}\chi_{1,j}^{2}\quad\text{and}\quad\hat{\sigma}_{n}^{2}\sum_{j=1}^{r}\hat{\Lambda}_{n,j}\chi_{1,j}^{2},$
respectively; thus $\hat{c}_{\alpha}$ is a natural estimator of $c_{\alpha}$. We reject $\mathcal{H}_{0}$ if $W_{n}>\hat{c}_{\alpha}$, equivalently with test function $1_{\{W_{n}>\hat{c}_{\alpha}\}}$. The following theorem shows that this test attains asymptotic level $\alpha$ under $\mathcal{H}_{0}$.
Theorem $3.$
Let Assumptions 14 hold. Under $\mathcal{H}_{0}$, $\mathsf{P}(W_{n}>\hat{c}_{\alpha})\to\alpha$ as $n\to\infty$.
Remark 6.
The critical value $\hat{c}_{\alpha}$ is obtained by Monte Carlo simulation, with a moment-matching approximation as a faster alternative.
4 Proofs
Throughout this section, all limits refer to the regime $n\to\infty$ with $p=p_{n}$ possibly diverging.
Proof of Lemma 2.1.
It suffices to show that $\hat{T}_{n}=T_{\gamma}$ under the event $\{\|\hat{\bm{\theta}}_{n}-\bm{\theta}_{0}\|_{\infty}\leq c_{n}\}$. For every $j\in\hat{T}_{n}^{c}$, we have $|\hat{\theta}_{nj}|\leq\tau_{n}$ and $|\theta_{0j}|-|\hat{\theta}_{nj}|\leq|\hat{\theta}_{nj}-\theta_{0j}|\leq c_{n}$. Then, we have $|\theta_{0j}|\leq c_{n}+\tau_{n}\leq\gamma$, which implies that $j\in T_{\gamma}^{c}$. Therefore, we obtain $\hat{T}_{n}^{c}\subset T_{\gamma}^{c}$, or equivalently, $T_{\gamma}\subset\hat{T}_{n}$. It holds that $T_{\gamma}^{c}=T_{W}\cup T_{0}^{c}$. For every $j\in T_{0}^{c}$, we have $|\hat{\theta}_{nj}|=|\hat{\theta}_{nj}-\theta_{0j}|\leq c_{n}\leq\tau_{n}$, which implies that $j\in\hat{T}_{n}^{c}$. Similarly, for every $j\in T_{W}$, it holds that $|\hat{\theta}_{nj}|\leq c_{n}+|\theta_{0j}|\leq c_{n}+\eta_{p}\leq\tau_{n}$, which implies that $j\in\hat{T}_{n}^{c}$. In conclusion, we have $T_{\gamma}^{c}\subset\hat{T}_{n}^{c}$. ∎
Proof of Lemma 2.2.
It follows from Bernstein’s inequality for sums of i.i.d. matrices established by Tropp (2012) that $\|\bm{J}_{nT_{\gamma},T_{\gamma}}-\bm{J}_{T_{\gamma},T_{\gamma}}\|_{\mathrm{op}}=o_{p}(1)$. ∎
Proof of Lemma 2.4.
(i) It holds that
$\mathsf{P}\left(\|n^{1/2}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}^{c}}\|_{2}^{2}>\zeta\right)\leq\,\mathsf{P}\left(\left\{n\|(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}^{c}}\|_{2}^{2}>\zeta\right\}\cap\{\hat{T}_{n}=T_{\gamma}\}\right)+\,\mathsf{P}(\hat{T}_{n}\neq T_{\gamma}).$
The bound
$n\|(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}^{c}}\|_{2}^{2}=n\sum_{j\in T_{\gamma}^{c}}|\theta_{0j}|^{2}\leq ns_{0}\eta_{p}^{2}$
holds on the event $\{\hat{T}_{n}=T_{\gamma}\}$ and hence,
$\mathsf{P}\left(\left\{n\|(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}^{c}}\|_{2}^{2}>\zeta\right\}\cap\{\hat{T}_{n}=T_{\gamma}\}\right)\leq\,\mathsf{P}\left(\left\{ns_{0}\eta_{n}^{2}>\zeta\right\}\cap\{\hat{T}_{n}=T_{\gamma}\}\right)\to 0.$
Moreover, Lemma 2.1 ensures $\mathsf{P}(\hat{T}_{n}\neq T_{\gamma})\to 0$.
[…]
$\psi_{nT}(\bm{\theta}_{T})=\frac{1}{n}\sum_{i=1}^{n}\bm{J}_{T,T}^{-1}\bm{Z}_{iT}(Y_{i}-\bm{\theta}_{T}^{\top}\bm{Z}_{iT})$
and then
$n^{1/2}\psi_{nT_{\gamma}}(\bm{\theta}_{0T_{\gamma}})1_{\{\hat{T}_{n}=T_{\gamma}\}}=\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{J}_{n\hat{T}_{n},\hat{T}_{n}}n^{1/2}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{\hat{T}_{n}}1_{\{\hat{T}_{n}=T_{\gamma}\}}.$
Since $\|\bm{J}_{nT_{\gamma},T_{\gamma}}-\bm{J}_{T_{\gamma},T_{\gamma}}\|_{\mathrm{op}}=o_{p}(1)$ and $1_{\{\hat{T}_{n}=T_{\gamma}\}}=1+o_{p}(1)$, we obtain
$n^{1/2}\bm{u}_{T_{\gamma}}^{\top}\psi_{nT_{\gamma}}(\bm{\theta}_{0T_{\gamma}})1_{\{\hat{T}_{n}=T_{\gamma}\}}=\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{J}_{nT_{\gamma},T_{\gamma}}n^{1/2}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}1_{\{\hat{T}_{n}=T_{\gamma}\}}$
and
$n^{-1/2}\sum_{i=1}^{n}\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{Z}_{iT_{\gamma}}\epsilon_{i}=\bm{u}_{T_{\gamma}}^{\top}{n}^{1/2}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}+o_{p}(1).$
Moreover,
$n^{-1}\sum_{i=1}^{n}\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{Z}_{iT_{\gamma}}\bm{Z}_{iT_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{u}_{T_{\gamma}}\epsilon_{i}^{2}\to^{p}\sigma^{2}\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{u}_{T_{\gamma}}.$
Furthermore, since $\bm{Z}_{1},\ldots,\bm{Z}_{n}$ and $\epsilon_{1},\ldots,\epsilon_{n}$ are i.i.d. copies of sub-Gaussian variables $\bm{Z}$ and $\epsilon$, respectively, and $s_{\gamma}$ is independent of $n$ and $p$, we have $\mathsf{E}[|\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{Z}_{T_{\gamma}}|^{2+\delta}]=O(1)$ and $\mathsf{E}[|\epsilon|^{2+\delta}]=O(1)$ for any $\delta>0$. Thus,
$\sum_{i=1}^{n}\,\mathsf{E}[|n^{-1/2}\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{Z}_{iT_{\gamma}}\epsilon_{i}|^{2+\delta}]\leq n^{-\delta/2}\,\mathsf{E}[|\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{Z}_{T_{\gamma}}|^{2+\delta}]\,\mathsf{E}[|\epsilon|^{2+\delta}]\\ =o(1).$
[…]
Proof of Corollary 3.1.
Let $\mathcal{H}_{0}$ hold. Define a linear operator $\mathcal{T}_{A}:\ell^{2}\to\ell^{2}$ by
$\langle\bm{e}_{j},\mathcal{T}_{A}(\bm{x})\rangle=\langle\bm{a}_{j},\bm{x}\rangle 1_{\{j\in[q]\}}\quad(j\in\mathbb{N})$
[…]
Proof of Lemma 3.2.
First, we show the consistency of $\hat{\sigma}_{n}^{2}$. Decompose $\hat{\sigma}_{n}^{2}$ as $\hat{\sigma}_{n}^{2}=S_{n1}-2S_{n2}+S_{n3}$, where
$S_{n1}=n^{-1}\sum_{i=1}^{n}\epsilon_{i}^{2},\quad S_{n2}=n^{-1}\sum_{i=1}^{n}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{\hat{T}_{n}}^{\top}\bm{Z}_{i\hat{T}_{n}}\epsilon_{i},\quad S_{n3}=n^{-1}\sum_{i=1}^{n}\left\{(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{\hat{T}_{n}}^{\top}\bm{Z}_{i\hat{T}_{n}}\right\}^{2}.$
By the law of large numbers, $S_{n1}\to^{p}\sigma_{0}^{2}$. To prove $S_{n2}=o_{p}(1)$, it suffices to show that
$\mathsf{P}\left(n^{-1}\left|\sum_{i=1}^{n}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{\hat{T}_{n}}^{\top}\bm{Z}_{i\hat{T}_{n}}\epsilon_{i}\right|>\zeta\right)\to 0,$
for any $\zeta>0$. As well as the proof of Lemma 2.4, we have
$\,\mathsf{P}\left(n^{-1}\left|\sum_{i=1}^{n}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{\hat{T}_{n}}^{\top}\bm{Z}_{i\hat{T}_{n}}\epsilon_{i}\right|>\zeta\right)$
$\leq\,\mathsf{P}\left(\left\{n^{-1}\left|\sum_{i=1}^{n}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}^{\top}\bm{Z}_{iT_{\gamma}}\epsilon_{i}\right|>\zeta\right\}\cap\{\hat{T}_{n}=T_{\gamma}\}\right)+\,\mathsf{P}(\hat{T}_{n}\neq T_{\gamma})$
$\leq\,\mathsf{P}\left(n^{-1}\left|\sum_{i=1}^{n}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}^{\top}\bm{Z}_{iT_{\gamma}}\epsilon_{i}\right|>\zeta\right)+\,\mathsf{P}(\hat{T}_{n}\neq T_{\gamma}).$
We thus get
$n^{-1}\left|\sum_{i=1}^{n}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}^{\top}\bm{Z}_{iT_{\gamma}}\epsilon_{i}\right|\leq\|(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}\|_{2}\left\|n^{-1}\sum_{i=1}^{n}\bm{Z}_{iT_{\gamma}}\epsilon_{i}\right\|_{2}=o_{p}(1)\cdot O_{p}(1)=o_{p}(1),$
because $n^{1/2}\|\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0}\|_{2}$ is $O_{p}(1)$ and $s_{\gamma}$ is independent of $n$ and $p$. Moreover, Lemma 2.1 implies $\mathsf{P}(\hat{T}_{n}\neq T_{\gamma})=o(1)$. To prove $S_{n3}=o_{p}(1)$, it suffices to show that
$\frac{1}{n}\sum_{i=1}^{n}\left\{(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}^{\top}\bm{Z}_{iT_{\gamma}}\right\}^{2}=o_{p}(1).$
It follows that
$n^{-1}\sum_{i=1}^{n}\left\{(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}^{\top}\bm{Z}_{iT_{\gamma}}\right\}^{2}\leq\|(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}\|_{2}^{2}\left\|n^{-1}\sum_{i=1}^{n}\bm{Z}_{i}\right\|_{2}^{2}=o_{p}(1)O_{p}(1)=o_{p}(1).$
Next, we show the consistency of $\hat{\Lambda}_{n,1},\ldots,\hat{\Lambda}_{n,r}$. It follows from Weyl’s inequality and Lemma 2.2 that $\max_{j\in[r]}|\hat{\Lambda}_{n,j}-\Lambda_{j}|\leq\|\hat{\bm{\Sigma}}_{nA}-\bm{\Sigma}_{A}\|_{\mathrm{op}}=o_{p}(1)$. ∎
Proof of Theorem $3.$
Let $\mathcal{H}_{0}$ hold. Since $\|\bm{R}_{A}\|_{2}^{2}$ has a continuous distribution function, we have $\hat{c}_{\alpha}\to^{p}c_{\alpha}$ by Lemma 3.2. Hence, by Slutsky’s lemma, $(W_{n},\hat{c}_{\alpha})\to^{d}(\|\bm{R}_{A}\|_{2}^{2},c_{\alpha})$. Since $\mathsf{P}(\|\bm{R}_{A}\|_{2}^{2}=c_{\alpha})=0$, we have $1_{\{W_{n}>\hat{c}_{\alpha}\}}\to^{d}1_{\{\|\bm{R}_{A}\|_{2}^{2}>c_{\alpha}\}}$ by the continuous mapping theorem. Finally, it follows from the the Portmanteau theorem that $\mathsf{P}(W_{n}>\hat{c}_{\alpha})\to\,\mathsf{P}(\|\bm{R}_{A}\|_{2}^{2}>c_{\alpha})=\alpha$. ∎
Acknowledgements
The second author was supported by Japan Society for the Promotion of Science KAKENHI Grant Number 25K07133.