### v3

# Asymptotically normal estimators in high-dimensional linear regression

Abstract
We establish asymptotic normality for estimators in high‑dimensional linear regression by proving weak convergence in a separable Hilbert space, thereby enabling direct use of standard asymptotic tools, for example, the continuous mapping theorem. The approach allows the number of non‑zero coefficients to grow, provided only a fixed number have moderate magnitude. As an application, we test linear hypotheses with a statistic whose null limit is a finite weighted sum of independent chi‑squared variables, yielding plug‑in critical values with asymptotically correct size.
keywords: high-dimensional inference; linear regression; sparsity; weak convergence
1 Introduction
We develop a new framework for establishing the asymptotic normality of estimators of high-dimensional parameters via weak convergence in the Hilbert space $\ell^{2}$, thereby enabling the use of standard asymptotic tools for broad classes of continuous functionals. Existing results typically assume fixed sparsity, namely that the number of non-zero coefficients does not grow, or restrict attention to low-dimensional linear functionals. We relax these restrictions and allow the sparsity level to diverge; in particular, the number of non-zero coefficients of small magnitude may increase with the sample size. For clarity of exposition, we focus on linear regression with independent and identically distributed (i.i.d.) sub-Gaussian covariates and errors.
Numerous methods for high-dimensional sparse estimation have been proposed, including the least absolute shrinkage and selection operator (lasso) (Tibshirani, 1996), the adaptive lasso (Zou, 2006), the Dantzig selector (Candes and Tao, 2007), and losses with nonconvex penalties such as the smoothly clipped absolute deviation and the minimax concave penalty (Fan and Li, 2001; Zhang, 2010). Foundational results include oracle inequalities and support-recovery guarantees (Bickel et al., 2009; Wainwright, 2009). On the inferential side, post-selection procedures have been analyzed (Belloni and Chernozhukov, 2013); separately, two-step estimators for stochastic process models have been developed (Fujimori and Tsukuda, 2026); and de-biasing methods yield asymptotically normal estimators of low-dimensional functionals (Zhang and Zhang, 2014; van de Geer et al., 2014; Javanmard and Montanari, 2014).
Accordingly, an extensive literature has developed non-asymptotic, high-dimensional central limit theorems and Gaussian approximations under metrics such as the Kolmogorov and Wasserstein distances. Representative results include Gaussian approximations for maxima of sums of centred independent random vectors (Chernozhukov et al., 2013, 2017), Stein kernel–based approximations accommodating exchangeable pairs and certain nonlinear statistics (Fang and Koike, 2021), rate refinements (Koike, 2023), and extensions to convex polytopes and degenerate cases (Fang et al., 2023).
Rather than deriving metric‑specific non-asymptotic bounds, we establish weak convergence of a high-dimensional estimator in $\ell^{2}$, bringing standard asymptotic tools such as the continuous mapping theorem to bear on a broad class of continuous functionals. Coordinatewise normality and normality of fixed‑dimensional linear functionals follow immediately. We illustrate the framework by constructing a test of linear hypotheses, contrasting it with projection‑based procedures that accommodate dense nuisance parameters (Zhu and Bradic, 2018) and hypothesis‑adaptive tests designed for high power (Zhang et al., 2025); our approach handles non-sparsity via a different route.
Throughout the paper, we use the following notation. We write $k]=\{1,\ldots,k\}$ for $k\in\mathbb{N}$. For $q\in\mathbb{N}\cup\{\infty\}$, write $\|\bm{v}\|_{q}$ for the $\ell^{q}$ norm: for $\bm{v}=(v_{1},\ldots,v_{p})^{\top}\in\mathbb{R}^{p}$ with $p\in\mathbb{N}$, $\|\bm{v}\|_{q}=(\sum_{j=1}^{p}|v_{j}|^{q})^{1/q}$ $(q<\infty)$, $\|\bm{v}\|_{\infty}=\max_{1\leq j\leq p}|v_{j}|$. For an infinite sequence $\bm{v}=(v_{j})_{j\geq 1}\in\ell^{q}$, the same formulae apply. For a bounded linear operator $\bm{A}:\ell^{2}\to\ell^{2}$ or a matrix $\bm{A}$ in finite dimensions, let $\|\bm{A}\|_{\mathrm{op}}=\sup_{\|\bm{x}\|_{2}=1}\|\bm{A}\bm{x}\|_{2}$ denote the $\ell^{2}$‑induced operator norm. For a vector $\bm{v}=(v_{j})_{j\geq 1}$ and an index set $T\subset\mathbb{N}$ with $|T|<\infty$, write $\bm{v}_{T}=(v_{j})_{j\in T}$, which is viewed as an element of $\mathbb{R}^{|T|}$. In the finite‑dimensional case $\bm{v}\in\mathbb{R}^{p}$, we restrict to $T\subset[p]$. Similarly, for a $p\times p$ matrix $\bm{A}=(A_{ij})_{i,j\in[p]}$ and index sets $T,T^{\prime}\subset[p]$, write $\bm{A}_{T,T^{\prime}}=(A_{ij})_{i\in T,j\in T^{\prime}}$. For a random variable $X$, we write the sub-Gaussian norm of $X$, if it exists, as $\|X\|_{\psi_{2}}$, that is, $\|X\|_{\psi_{2}}=\inf\{C>0:\,\mathsf{E}[\exp({X^{2}}/{C^{2}})]\leq 2\}$. For a $d$-dimensional random vector $\bm{X}$, the sub-Gaussian norm of $\bm{X}$ is $\|\bm{X}\|_{\psi_{2}}=\sup_{\|\bm{u}\|_{2}=1}\|\bm{u}^{\top}\bm{X}\|_{\psi_{2}}$. For $d\in\mathbb{N}$, $\bm{\mu}\in\mathbb{R}^{d}$, and $\bm{V}\in\mathbb{R}^{d\times d}$ positive semi‑definite, we write $\mathcal{N}_{d}(\bm{\mu},\bm{V})$ for the $d$-dimensional normal distribution with mean $\bm{\mu}$ and covariance $\bm{V}$, with the subscript omitted when $d=1$. For random elements $X$ and $Y$ and a sequence $\{X_{n}\}_{n\in\mathbb{N}}$ of random elements, we write $X\stackrel{{\scriptstyle d}}{{=}}Y$ if $X$ and $Y$ have identical distributions, and we write $X_{n}\to^{d}X$ and $X_{n}\to^{p}X$ for convergence in distribution and in probability as $n\to\infty$, respectively.
2 Model set‑up and main result
For each $n\in\mathbb{N}$, we work with the triangular array $\{(Y_{i,n},\bm{Z}_{i,n})\}_{i\in[n]}$ satisfying
$Y_{i,n}=\bm{\theta}_{n}^{\top}\bm{Z}_{i,n}+\epsilon_{i,n},\quad i\in[n],$
where $\bm{\theta}_{n}\in\mathbb{R}^{p_{n}}$ and $\bm{Z}_{i,n}=(Z_{i1,n},\ldots,Z_{ip_{n},n})^{\top}\in\mathbb{R}^{p_{n}}$ for $i\in[n]$. Assume that $\{(\bm{Z}_{i,n},\epsilon_{i,n})\}_{i\in[n]}$ are i.i.d. with $\mathsf{E}(\bm{Z}_{1,n})=\bm{0}$, $\mathsf{E}(\epsilon_{1,n})=0$, $\mathsf{V}(\epsilon_{1,n})=\sigma_{n}^{2}>0$, and $\bm{Z}_{1,n}$ independent of $\epsilon_{1,n}$. Let $\bm{\theta}_{0,n}$ denote the true parameter vector. Unless otherwise stated, throughout this paper all limits are taken as $n\to\infty$, allowing $p_{n}$ to diverge. Hereafter, we suppress the subscript $n$ when no confusion can arise. Write $\bm{\theta}_{0}=(\theta_{01},\ldots,\theta_{0p})^{\top}$. Define the index sets
$T_{0}:=\{j\in[p]:\theta_{0j}\neq 0\},\quad T_{\gamma}:=\{j\in[p]:|\theta_{0j}|>\gamma\},\quad T_{W}:=\{j\in[p]:\theta_{\min}\leq|\theta_{0j}|\leq\eta_{p}\},$
where $\theta_{\min}:=\min_{j\in T_{0}}|\theta_{0j}|$, $\gamma>0$ is a fixed constant independent of $n$ and $p$, and $\eta_{p}$ is a positive sequence depending on $p$. We write $s_{0}:=|T_{0}|$ for the support size, which may diverge, and $s_{\gamma}:=|T_{\gamma}|$ for the number of coefficients of moderate magnitude.
Assumption 1.
The following conditions hold:
  1. (i)
    $T_{\gamma}$ is fixed and independent of $n$ and $p$.
  2. (ii)
    $T_{0}\setminus T_{\gamma}=T_{W}$.
  3. (iii)
    $ns_{0}\eta_{p}^{2}\to 0$.
  4. (iv)
    There exists $\theta_{\max}>0$, independent of $n$ and $p$, such that $\|\bm{\theta}_{0,n}\|_{\infty}\leq\theta_{\max}$.
  5. (v)
    $\sigma^{2}$ is fixed and independent of $n$ and $p$.
Remark 1.
Assumption 1-(ii) states that the non-zero coefficients outside $T_{\gamma}$ are precisely the weak signals, and Assumption 1-(iii) requires that their aggregate magnitude is asymptotically negligible. Under Assumption 1, $(\bm{\theta}_{0,n}^{\top},0,0,\ldots)^{\top}\in\mathbb{R}^{\infty}$ lies in $\ell^{2}$ uniformly in $n$, because
$\|\bm{\theta}_{0,n}\|_{2}^{2}=\sum_{j\in T_{\gamma}}|\theta_{0j}|^{2}+\sum_{j\in T_{0}\setminus T_{\gamma}}|\theta_{0j}|^{2}\leq s_{\gamma}\theta_{\max}^{2}+s_{0}\eta_{p}^{2}=O(1).$
3
Let $\hat{\bm{\theta}}_{n}=(\hat{\theta}_{n1},\ldots,\hat{\theta}_{np})^{\top}$ be a sparse estimator for $\bm{\theta}$ such that $\mathsf{P}(\|\hat{\bm{\theta}}_{n}-\bm{\theta}_{0}\|_{\infty}>r_{n})\to 0$ for some rate $r_{n}\to 0$. Given a threshold sequence $\{\tau_{n}\}_{n\in\mathbb{N}}$, define the sequence $\{\hat{T}_{n}\}_{n\in\mathbb{N}}$ of an estimator of $T_{0}$ by
$\hat{T}_{n}=\{j\in[p]:|\hat{\theta}_{nj}|>\tau_{n}\}.$
Remark 2.
Under standard conditions, the lasso estimator and the Dantzig selector obey the high-probability bound $\|\hat{\bm{\theta}}_{n}-\bm{\theta}_{0}\|_{\infty}\leq Cn^{-1/2}(\log{p})^{1/2}$ for some $C>0$. Accordingly, in such cases we may take $r_{n}\asymp n^{-1/2}(\log p)^{1/2}$.
The following condition suffices to ensure selection consistency.
Assumption 2.
The following condition holds: $r_{n}+\eta_{p}<\tau_{n}<\gamma-r_{n}$.
Lemma 2.1.
Let Assumptions 12 hold. Then, $\mathsf{P}(\hat{T}_{n}=T_{\gamma})\to 1$.
The post-selection ordinary least squares estimator $\tilde{\bm{\theta}}_{n}=(\tilde{\theta}_{n1},\ldots,\tilde{\theta}_{np})^{\top}$ is defined as the solution to the estimating equation
$n^{-1}\sum_{i=1}^{n}\bm{Z}_{i\hat{T}_{n}}(Y_{i}-\bm{\theta}_{\hat{T}_{n}}^{\top}\bm{Z}_{i\hat{T}_{n}})=\bm{0},\quad\bm{\theta}_{\hat{T}_{n}^{c}}=\bm{0}.$
[…]
where $\{\bm{e}_{j}\}_{j\in\mathbb{N}}$ is the canonical basis of $\mathbb{R}^{\infty}$. Define
$\hat{\bm{J}}_{n}:=n^{-1}\sum_{i=1}^{n}\bm{Z}_{i}\bm{Z}_{i}^{\top},\quad\bm{J}_{n}:=\,\mathsf{E}[\bm{Z}_{1}\bm{Z}_{1}^{\top}].$
We suppose that there exists a positive definite matrix $\bm{J}_{T_{\gamma},T_{\gamma}}=\lim\bm{J}_{nT_{\gamma},T_{\gamma}}$. In addition, we impose the following conditions.
Assumption 3.
The following conditions hold:
  1. (i)
    There exists $\lambda_{*}$, independent of $n$ and $p$, such that $0<\lambda_{*}\leq\Lambda_{\min}(\bm{J}_{nT_{\gamma},T_{\gamma}})$.
  2. (ii)
    $\sup_{n\in\mathbb{N}}\max_{j\in[p_{n}]}\|Z_{1j,n}\|_{\psi_{2}}<\infty$.
  3. (iii)
    $\sup_{n\in\mathbb{N}}\|\epsilon_{1,n}\|_{\psi_{2}}<\infty$.
Lemma 2.2.
Let Assumptions 1 and 3 hold. Then, $\|\hat{\bm{J}}_{nT_{\gamma},T_{\gamma}}-\bm{J}_{nT_{\gamma},T_{\gamma}}\|_{\mathrm{op}}=o_{p}(1)$.
We now state the main limit theorem.
Theorem 2.3.
[…]
Remark 3.
Fujimori and Tsukuda (2026) analyze asymptotic normality for high‑dimensional estimators via weak convergence in $\ell^{2}$, including settings with a possibly infinite‑dimensional nuisance parameter and dependent observations. Their analysis is confined to the fixed-sparsity regime, which materially simplifies the arguments. In contrast, we allow $s_{0}\to\infty$ while keeping $T_{\gamma}$ fixed.
3 Application: testing linear hypotheses
As a direct application of Theorem 2.3, we consider testing linear hypotheses for $\bm{\theta}$; the resulting test statistic and its null limit are stated below. Let $q(\leq p)$ be an integer, $\bm{A}\in\mathbb{R}^{q\times p}$ a given matrix, $\bm{b}\in\mathbb{R}^{q}$ a given vector. Let $\bm{\Sigma}_{A}=\bm{A}_{[q],T_{\gamma}}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{A}_{[q],T_{\gamma}}^{\top}$. Moreover, let $r$ be the rank of $\bm{\Sigma}_{A}$, and let $\Lambda_{1}\leq\ldots\leq\Lambda_{r}$ be the non-zero eigenvalues of $\bm{\Sigma}_{A}$. Since $s_{\gamma}$ is independent of $n$ and $p$, we have $r\leq\min(s_{\gamma},q)$, so that $r$ is fixed. To test $\mathrm{H}_{0}:\bm{A}\bm{\theta}=\bm{b}$ against $\mathrm{H}_{1}:\bm{A}\bm{\theta}\neq\bm{b}$, consider
$W_{n}:=n\|\bm{A}\tilde{\bm{\theta}}_{n}-\bm{b}\|_{2}^{2}.$
We now state the condition on $q$.
Assumption 4.
The following condition holds: $q$ is fixed and independent of $n$ and $p$.
The asymptotic null distribution of $W_{n}$ is given in the following corollary.
Corollary 3.1.
Let Assumptions 14 hold. Under $\mathrm{H}_{0}$,
$W_{n}\to^{d}\sigma^{2}\sum_{j=1}^{r}\Lambda_{j}\chi_{1,j}^{2},$
where $\chi_{1,1}^{2},\ldots,\chi_{1,r}^{2}$ are independent and $\chi^{2}(1)$-distributed random variables.
We consider the following simple estimators of $\sigma^{2}$ and $\Lambda_{1},\ldots,\Lambda_{r}$. For $\sigma^{2}$, define
$\hat{\sigma}_{n}^{2}=n^{-1}\sum_{i=1}^{n}\left(Y_{i}-\tilde{\bm{\theta}}_{n}^{\top}\bm{Z}_{i}\right)^{2}.$
For $\Lambda_{j}$ ($j\in[r]$), define $\hat{\Lambda}_{n,j}$ as the $j$th largest eigenvalue of
$\hat{\bm{\Sigma}}_{nA}=\bm{A}_{[q],T_{\gamma}}\hat{\bm{J}}_{nT_{\gamma},T_{\gamma}}^{-1}\bm{A}_{[q],T_{\gamma}}^{\top}.$
The following lemma establishes consistency of these estimators.
Lemma 3.2.
Let Assumptions 14 hold. Then, $\hat{\sigma}_{n}^{2}\to^{p}\sigma^{2}$ and $\max_{j\in[r]}|\hat{\Lambda}_{n,j}-\Lambda_{j}|\to^{p}0$.
Let $c_{\alpha}$ and $\hat{c}_{\alpha}$ denote the upper $\alpha$-quantiles of the distributions of
$\sigma^{2}\sum_{j=1}^{r}\Lambda_{j}\chi_{1,j}^{2}\quad\text{and}\quad\hat{\sigma}_{n}^{2}\sum_{j=1}^{r}\hat{\Lambda}_{n,j}\chi_{1,j}^{2},$
respectively; thus $\hat{c}_{\alpha}$ is a natural estimator of $c_{\alpha}$. We reject $\mathrm{H}_{0}$ if $W_{n}>\hat{c}_{\alpha}$, equivalently with test function $\varphi_{n}=1_{\{W_{n}>\hat{c}_{\alpha}\}}$. The following theorem shows that this test attains asymptotic level $\alpha$ under $\mathrm{H}_{0}$.
Theorem $3.$
Let Assumptions 14 hold. Under $\mathrm{H}_{0}$, $\mathsf{P}(\varphi_{n}=1)\to\alpha$.
Remark 4.
The critical value $\hat{c}_{\alpha}$ is obtained by Monte Carlo simulation, with a moment-matching approximation as a faster alternative.
Remark 5.
Zhu and Bradic (2018) and Zhang et al. (2025) conduct statistical inference with dense $\bm{\theta}$ by imposing sparsity on orthogonalized loading directions obtained after suitable projection. By contrast, we impose no sparsity on $\bm{A}$ and instead proceed under stronger assumptions on $\bm{\theta}$, thereby situating our results in a different inferential setting.
Remark 6.
Although we assume $q$ is fixed, the results can be extended to the regime $q=q_{n}\to\infty$,provided that $\bm{A}$ defines a bounded linear operator on $\ell^{2}$.
4 Proofs
Proof of Lemma 2.1.
It suffices to show that $\hat{T}_{n}=T_{\gamma}$ on the event $\{\|\hat{\bm{\theta}}_{n}-\bm{\theta}_{0}\|_{\infty}\leq r_{n}\}$. For every $j\in\hat{T}_{n}^{c}$, we have $|\hat{\theta}_{nj}|\leq\tau_{n}$ and $|\theta_{0j}|-|\hat{\theta}_{nj}|\leq|\hat{\theta}_{nj}-\theta_{0j}|\leq r_{n}$. Then, we have $|\theta_{0j}|\leq r_{n}+\tau_{n}\leq\gamma$, which implies that $j\in T_{\gamma}^{c}$. Therefore, we obtain $\hat{T}_{n}^{c}\subset T_{\gamma}^{c}$, or equivalently, $T_{\gamma}\subset\hat{T}_{n}$. It holds that $T_{\gamma}^{c}=T_{W}\cup T_{0}^{c}$. For every $j\in T_{0}^{c}$, we have $|\hat{\theta}_{nj}|=|\hat{\theta}_{nj}-\theta_{0j}|\leq r_{n}\leq\tau_{n}$, which implies that $j\in\hat{T}_{n}^{c}$. Similarly, for every $j\in T_{W}$, it holds that $|\hat{\theta}_{nj}|\leq r_{n}+|\theta_{0j}|\leq r_{n}+\eta_{p}\leq\tau_{n}$, which implies that $j\in\hat{T}_{n}^{c}$. In conclusion, we have $T_{\gamma}^{c}\subset\hat{T}_{n}^{c}$. ∎
Proof of Lemma 2.2.
The conclusion follows from the law of large numbers for triangular arrays. ∎
Proof of Theorem 2.3.
It suffices to establish the following (i) and (ii); see, for example, Chapter 1.8 of van der Vaart and Wellner (2023): (i) For every $\zeta>0$, $\mathsf{P}(\|n^{1/2}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}^{c}}\|_{2}^{2}>\zeta)\to 0$. (ii) For every $\bm{u}\in\ell^{2}$, $𝓡\langle\bm{u},\bm{\mathcal{R}}_{n}\rangle\to^{d}\langle\bm{u},\bm{\mathcal{R}}\rangle$.
(i) It holds that
$\mathsf{P}\left(\|n^{1/2}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}^{c}}\|_{2}^{2}>\zeta\right)\leq\,\mathsf{P}\left(\left\{n\|(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}^{c}}\|_{2}^{2}>\zeta\right\}\cap\{\hat{T}_{n}=T_{\gamma}\}\right)+\,\mathsf{P}(\hat{T}_{n}\neq T_{\gamma}).$
The bound
$n\|(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}^{c}}\|_{2}^{2}=n\sum_{j\in T_{\gamma}^{c}}|\theta_{0j}|^{2}\leq ns_{0}\eta_{p}^{2}$
holds on the event $\{\hat{T}_{n}=T_{\gamma}\}$ and hence,
$\mathsf{P}\left(\left\{n\|(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}^{c}}\|_{2}^{2}>\zeta\right\}\cap\{\hat{T}_{n}=T_{\gamma}\}\right)\leq\,\mathsf{P}\left(\left\{ns_{0}\eta_{n}^{2}>\zeta\right\}\cap\{\hat{T}_{n}=T_{\gamma}\}\right)\to 0.$
Moreover, Lemma 2.1 ensures $\mathsf{P}(\hat{T}_{n}\neq T_{\gamma})\to 0$.
[…]
$\psi_{nT}(\bm{\theta}_{T})=n^{-1}\sum_{i=1}^{n}\bm{J}_{T,T}^{-1}\bm{Z}_{iT}(Y_{i}-\bm{\theta}_{T}^{\top}\bm{Z}_{iT})$
and then
$n^{1/2}\psi_{nT_{\gamma}}(\bm{\theta}_{0T_{\gamma}})1_{\{\hat{T}_{n}=T_{\gamma}\}}=\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\hat{\bm{J}}_{n\hat{T}_{n},\hat{T}_{n}}n^{1/2}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{\hat{T}_{n}}1_{\{\hat{T}_{n}=T_{\gamma}\}}.$
Since $\|\hat{\bm{J}}_{nT_{\gamma},T_{\gamma}}-\bm{J}_{T_{\gamma},T_{\gamma}}\|_{\mathrm{op}}\leq\|\hat{\bm{J}}_{nT_{\gamma},T_{\gamma}}-\bm{J}_{nT_{\gamma},T_{\gamma}}\|_{\mathrm{op}}+\|\bm{J}_{nT_{\gamma},T_{\gamma}}-\bm{J}_{T_{\gamma},T_{\gamma}}\|_{\mathrm{op}}=o_{p}(1)$ and $1_{\{\hat{T}_{n}=T_{\gamma}\}}=1+o_{p}(1)$, we obtain
$n^{1/2}\bm{u}_{T_{\gamma}}^{\top}\psi_{nT_{\gamma}}(\bm{\theta}_{0T_{\gamma}})1_{\{\hat{T}_{n}=T_{\gamma}\}}=\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\hat{\bm{J}}_{nT_{\gamma},T_{\gamma}}n^{1/2}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}1_{\{\hat{T}_{n}=T_{\gamma}\}}$
and
$n^{-1/2}\sum_{i=1}^{n}\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{Z}_{iT_{\gamma}}\epsilon_{i}=\bm{u}_{T_{\gamma}}^{\top}{n}^{1/2}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}+o_{p}(1).$
Moreover, the sub-Gaussian bound implies
$n^{-1}\sum_{i=1}^{n}\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{Z}_{iT_{\gamma}}\bm{Z}_{iT_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{u}_{T_{\gamma}}\epsilon_{i}^{2}\to^{p}\sigma^{2}\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{u}_{T_{\gamma}},$
because
$n^{-1}\sum_{i=1}^{n}\,\mathsf{E}[\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{Z}_{iT_{\gamma}}\bm{Z}_{iT_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{u}_{T_{\gamma}}\epsilon_{i}^{2}]=\sigma^{2}\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{J}_{nT_{\gamma},T_{\gamma}}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{u}_{T_{\gamma}}\to\sigma^{2}\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{u}_{T_{\gamma}}.$
Furthermore, since $s_{\gamma}$ is independent of $n$ and $p$, the sub-Gaussian bound implies $\mathsf{E}[|\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{Z}_{1T_{\gamma}}|^{2+\delta}]=O(1)$ and $\mathsf{E}[|\epsilon_{1}|^{2+\delta}]=O(1)$ for any $\delta>0$. Thus,
$\sum_{i=1}^{n}\,\mathsf{E}[|n^{-1/2}\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{Z}_{iT_{\gamma}}\epsilon_{i}|^{2+\delta}]\leq n^{-\delta/2}\,\mathsf{E}[|\bm{u}_{T_{\gamma}}^{\top}\bm{J}_{T_{\gamma},T_{\gamma}}^{-1}\bm{Z}_{1T_{\gamma}}|^{2+\delta}]\,\mathsf{E}[|\epsilon_{1}|^{2+\delta}]\\ =o(1).$
[…]
Proof of Corollary 3.1.
Let $\mathrm{H}_{0}$ hold. Define a linear operator $\mathcal{T}_{A}:\ell^{2}\to\ell^{2}$ by
$\langle\bm{e}_{j},\mathcal{T}_{A}(\bm{x})\rangle=\begin{cases}\langle\bm{a}_{j},\bm{x}\rangle&(j\in[q])\\ 0&(j\in\mathbb{N}\setminus[q])\end{cases}$
[…]
Proof of Lemma 3.2.
First, we show $\hat{\sigma}_{n}^{2}\to^{p}\sigma^{2}$. Decompose $\hat{\sigma}_{n}^{2}$ as $\hat{\sigma}_{n}^{2}=S_{n1}-2S_{n2}+S_{n3}$, where
$S_{n1}=n^{-1}\sum_{i=1}^{n}\epsilon_{i}^{2},\quad S_{n2}=n^{-1}\sum_{i=1}^{n}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{\hat{T}_{n}}^{\top}\bm{Z}_{i\hat{T}_{n}}\epsilon_{i},\quad S_{n3}=n^{-1}\sum_{i=1}^{n}\left\{(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{\hat{T}_{n}}^{\top}\bm{Z}_{i\hat{T}_{n}}\right\}^{2}.$
By the law of large numbers, $S_{n1}\to^{p}\sigma^{2}$. To prove $S_{n2}=o_{p}(1)$, it suffices to show that
$\mathsf{P}\left(n^{-1}\left|\sum_{i=1}^{n}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{\hat{T}_{n}}^{\top}\bm{Z}_{i\hat{T}_{n}}\epsilon_{i}\right|>\zeta\right)\to 0,$
for any $\zeta>0$. As well as the proof of Theorem 2.3, the left-hand side admits the bound
$\,\mathsf{P}\left(\left\{n^{-1}\left|\sum_{i=1}^{n}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}^{\top}\bm{Z}_{iT_{\gamma}}\epsilon_{i}\right|>\zeta\right\}\cap\{\hat{T}_{n}=T_{\gamma}\}\right)+\,\mathsf{P}(\hat{T}_{n}\neq T_{\gamma})$
$\leq\,\mathsf{P}\left(n^{-1}\left|\sum_{i=1}^{n}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}^{\top}\bm{Z}_{iT_{\gamma}}\epsilon_{i}\right|>\zeta\right)+\,\mathsf{P}(\hat{T}_{n}\neq T_{\gamma}).$
We have
$n^{-1}\left|\sum_{i=1}^{n}(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}^{\top}\bm{Z}_{iT_{\gamma}}\epsilon_{i}\right|\leq\|(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}\|_{2}\left\|n^{-1}\sum_{i=1}^{n}\bm{Z}_{iT_{\gamma}}\epsilon_{i}\right\|_{2}=o_{p}(1)O_{p}(1)=o_{p}(1),$
because $n^{1/2}\|\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0}\|_{2}=O_{p}(1)$ and $s_{\gamma}$ is independent of $n$ and $p$. Moreover, Lemma 2.1 implies $\mathsf{P}(\hat{T}_{n}\neq T_{\gamma})=o(1)$. To prove $S_{n3}=o_{p}(1)$, it suffices to show that
$n^{-1}\sum_{i=1}^{n}\left\{(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}^{\top}\bm{Z}_{iT_{\gamma}}\right\}^{2}=o_{p}(1).$
It follows that
$n^{-1}\sum_{i=1}^{n}\left\{(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}^{\top}\bm{Z}_{iT_{\gamma}}\right\}^{2}\leq\|(\tilde{\bm{\theta}}_{n}-\bm{\theta}_{0})_{T_{\gamma}}\|_{2}^{2}n^{-1}\sum_{i=1}^{n}\|\bm{Z}_{iT_{\gamma}}\|_{2}^{2}=o_{p}(1)O_{p}(1)=o_{p}(1).$
Next, from Weyl’s inequality and Lemma 2.2, we have
$\max_{j\in[r]}|\hat{\Lambda}_{n,j}-\Lambda_{j}|\leq\|\hat{\bm{\Sigma}}_{nA}-\bm{\Sigma}_{A}\|_{\mathrm{op}}\leq\|\hat{\bm{\Sigma}}_{nA}-\bm{\Sigma}_{nA}\|_{\mathrm{op}}+\|\bm{\Sigma}_{nA}-\bm{\Sigma}_{A}\|_{\mathrm{op}}=o_{p}(1),$
where $\bm{\Sigma}_{nA}=\bm{A}_{[q],T_{\gamma}}\bm{J}_{nT_{\gamma},T_{\gamma}}^{-1}\bm{A}_{[q],T_{\gamma}}^{\top}$. ∎
Proof of Theorem $3.$
Let $\mathrm{H}_{0}$ hold. Since $\|\bm{R}_{A}\|_{2}^{2}$ has a continuous distribution function, we have $\hat{c}_{\alpha}\to^{p}c_{\alpha}$ by Lemma 3.2. Hence, by Slutsky’s lemma, $(W_{n},\hat{c}_{\alpha})\to^{d}(\|\bm{R}_{A}\|_{2}^{2},c_{\alpha})$. Since $\mathsf{P}(\|\bm{R}_{A}\|_{2}^{2}=c_{\alpha})=0$, we have $\varphi_{n}\to^{d}1_{\{\|\bm{R}_{A}\|_{2}^{2}>c_{\alpha}\}}$ by the continuous mapping theorem. Finally, it follows from the the Portmanteau theorem that $\mathsf{P}(\varphi_{n}=1)\to\,\mathsf{P}(\|\bm{R}_{A}\|_{2}^{2}>c_{\alpha})=\alpha$. ∎
Acknowledgements
The second author was supported by the Japan Society for the Promotion of Science KAKENHI Grant Number 25K07133.